OpenStudy (habbababba1):

will fan and medal......The range of the function f(x) = x + 5 is {7, 9}. What is the function’s domain? {2, 4} {-2, -4} {12, 14} {-12, -14} {0, 5}

OpenStudy (fluffynose):

the range is the y value we can rewrite the given function f(x)=x+5 to y=x+5 and our given range are y=7 and y=9 so sub in ur y values into the equation to find the x value which are the domain... y=x+5

OpenStudy (jjcoolja12):

Let's solve for x. y=x+5 Step 1: Flip the equation. x+5=y Step 2: Add -5 to both sides. x+5+−5=y+−5 x=y−5 Answer: x=y−5 so its really x=y-5

OpenStudy (er.mohd.amir):

here range is value of f(x) and domain is value of x so find value of x for given f(x).

OpenStudy (shaik0124):

{2,4}

OpenStudy (anonymous):

id say 2,4 tbh

OpenStudy (jameshorton):

f(x) = 1/4x + 3 when the range is {3, -17, 27}? x = 0, -80, 120 Domain: {0,80,120} f(x) = -3x - 5 when the range is {-2, 0, 3}? x = -1, -5/3, -8/3